OpenStudy (anonymous):

For a book lying on the table, Why is the sum of normal reaction force and the book's weight not equal to zero?

OpenStudy (anonymous):

Newtons third law states that for every action there is an equal and opposite reaction. For a book lying on a table, the weight is is due to gravity acting upon its mass, such that \(F_b=mg\). The normal reaction is equal but opposite, and hence it is \(F_R=-mg\) the negative sign showing it is in the opposite direction. The sum of the forces is then \[F_b+F_R=mg+(-mg)=mg-mg=0\].